### v2

@@ Appendix A Algorithm Details
Algorithm 2 Expert-specific summation
[…]
  Initialize: Output tokens $y$ with shape $(E,D)$ initialized with 0.
  Parallel for $i=0$ to $E-1$ do
   Parallel for $j$ in range $(0,D,\texttt{BLK})\ \textbf{do}$
     $exp=\mathcal{R}[v[idx[i]]]$
    Initialize zero tensor $c$ with shape $(1,\texttt{BLK})$
[…]
    for $k$ in range $(idx[i],idx[i+1],\texttt{BLK})\ \textbf{do}$
     Parallel for $t=0$ to BLK  do
      load $x_{sub}[t]=x[v[k+t],j:j+\texttt{BLK}]$
     end for
      $c=c+\sum\limits_{t=0}^{\texttt{BLK}-1}x_{sub}[t]$
    end for
   Write back: $y[exp,j:j+\texttt{BLK}]=c$
   end for
  end for
  Output: Tensor $y$
@@
Appendix B Experimental Details
@@
forward ESMM $\bm{x}$ $(N,D_{1})$ $\bm{y}$ $(N,D_{2})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
$\bm{w}$ $(E,D_{1},D_{2})$
$\bm{b}$ $(E,D_{2})$
backward ESMM $\bm{x}$ $(N,D_{2})$ $\bm{y}$ $(N,D_{1})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
$\bm{w}$ $(E,D_{2},D_{1})$
ESS $\bm{x}$ $(N,D_{2})$ $\bm{y}$ $(E,D_{2})$ (WARP, TIMES) ($E$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
ESTMM $\bm{x}_{1}$ $(N,D_{1})$ $\bm{w}$ $(E,D_{1},D_{2})$ (WARP, TIMES) ($\lceil D_{2}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $E$)
$\bm{x}_{2}$ $(N,D_{2})$
ESFK $\bm{x}_{1}$ $(N,D_{1})$ $\bm{y}_{1}$ $(N,D_{1})$ (WARP, TIMES) ($E$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $\lceil N^{\prime}/\texttt{BLK}\rceil+\lceil D_{2}/\texttt{BLK}\rceil+$ $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
$\bm{x}_{2}$ $(N,D_{2})$ $\bm{y}_{2}$ $(E,D_{2})$
$\bm{w}_{1}$ $(E,D_{2},D_{1})$ $\bm{w}_{2}$ $(E,D_{1},D_{2})$
We also provide the Python code for the proxy task we used to examine the computing capacity of the heterogeneous devices. We adopt a for loop composed of large matrix multiplications with the same scale as the test program, as provided below:
import torch
import time
device = ’cuda’
size = 2048
times = 1024
start_time = time.time()
for j in range(times):
mat1 = torch.randn(size, size, device=device)
mat2 = torch.randn(size, size, device=device)
y = torch.matmul(mat1, mat2)
end_time = time.time()
print(end_time - start_time)
Appendix C Experimental Results
We provide the exact values for both memory usage analysis and average latency analysis in Table 7 and 8, respectively. Specifically, for latency analysis, we provide the average value for each case with 0.5k, 1k, 1.5k and 2k total steps.
We also visualize the validation accuracy of our method and Tutel under the same training time in Figure 25. We take Swin-MoE-Base with 4 experts and top-$1$ routing as the backbone network, training it on the homogeneous machine with the same batch size using Imagenet-1k dataset Deng et al. (2009). Since our HEXA-MoE can accelerate training process, its validation accuracy is much higher than baseline.
Table 7: Memory analysis with Tutel, MegaBlocks and HEXA-MoE on Swin-Transformer-MoE benchmark (Base and Small). Experiments are conducted on 2 homogeneous GPUs with automatic mixed precision in PyTorch and batch size 40 for all the experiments. We set the number of global experts to 8, and record the average GPU memory consumption (GB) on each device.
@@ Appendix C Experimental Results
Figure 25: Comparison on validation accuracy for Tutel and HEXA-MoE with the same training time. HEXA-MoE has a significant advantage over Tutel due to the highly improved efficiency on computation.

### v4

@@ Appendix B Algorithm Details
end for
end for
Output: Tensor $y$
Algorithm 3 Expert-specific summation
[…]
Initialize: Output tokens $y$ with shape $(E,D)$ initialized with 0.
Parallel for $i=0$ to $E-1$ do
Parallel for $j$ in range $(0,D,\texttt{BLK})\ \textbf{do}$
   $exp=\mathcal{R}[v[idx[i]]]$
  Initialize zero tensor $c$ with shape $(1,\texttt{BLK})$
[…]
  for $k$ in range $(idx[i],idx[i+1],\texttt{BLK})\ \textbf{do}$
   Parallel for $t=0$ to BLK  do
    load $x_{sub}[t]=x[v[k+t],j:j+\texttt{BLK}]$
   end for
    $c=c+\sum\limits_{t=0}^{\texttt{BLK}-1}x_{sub}[t]$
  end for
Write back: $y[exp,j:j+\texttt{BLK}]=c$
@@
We present the algorithm details for re-index vector construction as well as the expert-specific operators, and taking top-1 routing as an example for illustration, shown in Algorithm 1, 2, 3 and 4. In Algorithm 1, we re-arrange the routing choice vector $\mathcal{R}$ into re-indexed token vector $v$, along with the token index starting vector $idx$, which satisfies $idx[0]=0$ and $idx[E]=N^{\prime}$. We provide the length and range for the vectors in Table 4.
Table 4: Explanation for the auxiliary vectors.
vector length range
$\mathcal{R}$ $N$ $0\leq\mathcal{R}[i]<E$ for $0\leq i<N$
$v$ $N^{\prime}$ $0\leq v[i]<N$ for $0\leq i<N^{\prime}$
$idx$ $1+E$ $0\leq idx[i]<N^{\prime}$ for $0\leq i<1+E$
In Algorithm 2, 3 and 4, we assume that the feature dimension for each tensor is all divisible by BLK. Notice that when accessing the re-index vector $v$, we may get value -1, since the workload for each expert is dynamic. In this case, we would skip this index, and remain zero for the temporary loading variable.
We also provide the formulation for top-$k$ routing with our expert-specific operators, and take a comparison with top-$1$ routing. We denote the routing choice for top-$k$ as $\{\mathcal{R}_{i}(\boldsymbol{x})\}_{i=0}^{k-1}$, and other symbols keep consistent with Figure 2. We present the comparison on formulations for MoE forward and backward propagation in Table 5.
Forward 1 1st MLP $\boldsymbol{y}_{1}=\textit{ESMM}(\boldsymbol{x},\boldsymbol{W}_{1},\boldsymbol{b}_{1},\mathcal{R}(\boldsymbol{x}))$$\{\boldsymbol{y}_{1}^{i}\}_{i=0}^{k-1}:\boldsymbol{y}_{1}^{i}=\textit{ESMM}(\boldsymbol{x},\boldsymbol{W}_{1},\boldsymbol{b}_{1},\mathcal{R}_{i}(\boldsymbol{x}))$
2 Activation $\boldsymbol{y}_{2}=\mathcal{F}(\boldsymbol{y}_{1})$$\{\boldsymbol{y}_{2}^{i}\}_{i=0}^{k-1}:\boldsymbol{y}_{2}^{i}=\mathcal{F}(\boldsymbol{y}_{1}^{i})$
3 2nd MLP $\boldsymbol{y}=\textit{ESMM}(\boldsymbol{y}_{2},\boldsymbol{W}_{2},\boldsymbol{b}_{2},\mathcal{R}(\boldsymbol{x}))$ $\boldsymbol{y}=\sum_{i=0}^{k-1}\textit{ESMM}(\boldsymbol{y}_{2}^{i},\boldsymbol{W}_{2},\boldsymbol{b}_{2},\mathcal{R}_{i}(\boldsymbol{x}))$
Backward 4 2nd MLP $\frac{\partial\ell}{\partial\boldsymbol{b}_{2}}=\textit{ESS}(\frac{\partial\ell}{\partial\boldsymbol{y}},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\partial\boldsymbol{b}_{2}}=\sum_{i=0}^{k-1}\textit{ESS}(\frac{\partial\ell}{\partial\boldsymbol{y}},\mathcal{R}_{i}(\boldsymbol{x}))$
5 $\frac{\partial\ell}{\partial\boldsymbol{W}_{1}}=\textit{ESTMM}(\boldsymbol{x},\frac{\partial\ell}{\partial\boldsymbol{y}_{1}},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\partial\boldsymbol{W}_{1}}=\sum_{i=0}^{k-1}\textit{ESTMM}(\boldsymbol{x},\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}},\mathcal{R}_{i}(\boldsymbol{x}))$
6 $\frac{\partial\ell}{\partial\boldsymbol{y}_{2}}=\textit{ESMM}(\frac{\partial\ell}{\partial\boldsymbol{y}},\boldsymbol{W}_{2}^{T},\textit{null},\mathcal{R}(\boldsymbol{x}))$$\Big{\{}\frac{\partial\ell}{\partial\boldsymbol{y}_{2}^{i}}\Big{\}}_{i=0}^{k-1}:\frac{\partial\ell}{\partial\boldsymbol{y}_{2}^{i}}=\textit{ESMM}(\frac{\partial\ell}{\partial\boldsymbol{y}},\boldsymbol{W}_{2}^{T},\textit{null},\mathcal{R}_{i}(\boldsymbol{x}))$
7 Activation $\frac{\partial\ell}{\partial\boldsymbol{y}_{1}}=\frac{\partial\ell}{\partial\boldsymbol{y}_{2}}\odot\mathcal{F}^{\prime}(\boldsymbol{y}_{1})$$\Big{\{}\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}}\Big{\}}_{i=0}^{k-1}:\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}}=\frac{\partial\ell}{\partial\boldsymbol{y}_{2}^{i}}\odot\mathcal{F}^{\prime}(\boldsymbol{y}_{1}^{i})$
8 1st MLP $\frac{\partial\ell}{\boldsymbol{b}_{1}}=\textit{ESS}(\frac{\partial\ell}{\partial\boldsymbol{y}_{1}},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\boldsymbol{b}_{1}}=\sum_{i=0}^{k-1}\textit{ESS}(\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}},\mathcal{R}_{i}(\boldsymbol{x}))$
9 $\frac{\partial\ell}{\partial\boldsymbol{W}_{1}}=\textit{ESTMM}(\boldsymbol{x},\frac{\partial\ell}{\partial\boldsymbol{y}_{1}},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\partial\boldsymbol{W}_{1}}=\sum_{i=0}^{k-1}\textit{ESTMM}(\boldsymbol{x},\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}},\mathcal{R}_{i}(\boldsymbol{x}))$
10 $\frac{\partial\ell}{\partial\boldsymbol{x}}=\textit{ESMM}(\frac{\partial\ell}{\partial\boldsymbol{y}_{1}},\boldsymbol{W}_{1}^{T},\textit{null},\mathcal{R}(\boldsymbol{x}))$ $\frac{\partial\ell}{\partial\boldsymbol{x}}=\sum_{i=0}^{k-1}\textit{ESMM}(\frac{\partial\ell}{\partial\boldsymbol{y}_{1}^{i}},\boldsymbol{W}_{1}^{T},\textit{null},\mathcal{R}_{i}(\boldsymbol{x}))$
@@
forward ESMM $\boldsymbol{x}$ $(N,D_{1})$ $\boldsymbol{y}$ $(N,D_{2})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
$\boldsymbol{w}$ $(E,D_{1},D_{2})$
$\boldsymbol{b}$ $(E,D_{2})$
backward ESMM $\boldsymbol{x}$ $(N,D_{2})$ $\boldsymbol{y}$ $(N,D_{1})$ (WARP, TIMES) ($\lceil N^{\prime}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
$\boldsymbol{w}$ $(E,D_{2},D_{1})$
ESS $\boldsymbol{x}$ $(N,D_{2})$ $\boldsymbol{y}$ $(E,D_{2})$ (WARP, TIMES) ($E$, $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
ESTMM $\boldsymbol{x}_{1}$ $(N,D_{1})$ $\boldsymbol{w}$ $(E,D_{1},D_{2})$ (WARP, TIMES) ($\lceil D_{2}/\texttt{BLK}\rceil$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $E$)
$\boldsymbol{x}_{2}$ $(N,D_{2})$
ESFK $\boldsymbol{x}_{1}$ $(N,D_{1})$ $\boldsymbol{y}_{1}$ $(N,D_{1})$ (WARP, TIMES) ($E$, $\lceil D_{1}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$, $\lceil N^{\prime}/\texttt{BLK}\rceil+\lceil D_{2}/\texttt{BLK}\rceil+$ $\lceil D_{2}/(\texttt{TIMES}\cdot\texttt{BLK})\rceil$)
$\boldsymbol{x}_{2}$ $(N,D_{2})$ $\boldsymbol{y}_{2}$ $(E,D_{2})$
$\boldsymbol{w}_{1}$ $(E,D_{2},D_{1})$ $\boldsymbol{w}_{2}$ $(E,D_{1},D_{2})$
Appendix C Experimental Details
We also provide the PyTorch-style pseudocode for the proxy task we used to examine the computing capacity of the heterogeneous devices, as shown in Algorithm 5. We adopt a for loop composed of large matrix multiplications with the same scale as the test program.
Algorithm 5 PyTorch-style pseudocode of MoE pipeline.
1 import torch
2 import time
3
4 device = ’cuda’
5 size = 2048
6 times = 1024
7
8 start_time = time.time()
9 for j in range(times):
10 mat1 = torch.randn(size, size, device=device)
11 mat2 = torch.randn(size, size, device=device)
12 y = torch.matmul(mat1, mat2)
13 end_time = time.time()
14
15 print(end_time - start_time)
Appendix D Experimental Results
We provide the exact values for both memory footprint analysis and average latency analysis in Table 7 and 8, respectively. Specifically, for latency analysis, we provide the average value for each case with 0.5k, 1k, 1.5k and 2k total steps.
Table 7: Memory analysis with Tutel, MegaBlocks and HEXA-MoE on Swin-Transformer-MoE benchmark (Base and Small). Experiments are conducted on 2 homogeneous GPUs with automatic mixed precision in PyTorch and batch size 40 for all the experiments. We set the number of global experts to 8, and record the average GPU memory footprint (GB) on each device.